OpenStudy (anonymous):

Consider f(x)=sin^2(x)-cos(x) on the interval 0 is less than or equal to x which is less than or equal to pi. Determine the local extrema of f(x).

OpenStudy (anonymous):

\[f(\theta)=\sin^{2}(\theta)-\cos(\theta)\] \[0 \le \theta \le \pi\]

OpenStudy (anonymous):

I know that one of the local extrema occurs at x=0, but I am not sure of how to find the other two.